OpenStudy (anonymous):

If (x + 1)(x - 3) = 5, then which of the following statements is true?

OpenStudy (anonymous):

(x+1)(x-3)=5 =>x^2-3x+x-3=5 =>x^2-2x=5+3 =>x^2-2x-8=0 =>x^2-4x+2x-8=0 =>x(x-4)+2(x-4)=0 =>(x-4)(x+2)=0 so x=4,-2

OpenStudy (anonymous):

x + 1 = 0 or x - 3 = 0 x + 1 = 5 or x - 3 = 5 x - 4 = 0 or x + 2 = 0 these are the answer choices dow

OpenStudy (anonymous):

the answer is x-4=0 or x+2=0